OpenStudy (anonymous):

The number 320 million written in scientific notation would be _____. A.320 × 106. B.3.2 × 108. C.3.2 × 10-8. D.32 × 10-7.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 320,000,000 }\) \(\Large \color{MidnightBlue}{\Rightarrow 32 * 1,000,000 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3.2 * 10,000,000 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3.2 * 10^7 }\)

Parth (parthkohli):

Always do scientific notations like this.

OpenStudy (anonymous):

so whats the answer

Parth (parthkohli):

D. is the most appropriate.

OpenStudy (anonymous):

thx

Parth (parthkohli):

No, wait.

OpenStudy (apoorvk):

No, that would be 3.2 x 10^8

OpenStudy (anonymous):

Which of the following expressions is equal to 3 5 · 3 2? A.37. B.310. C.910. D.33.

Parth (parthkohli):

3.2 * 10^8 will be the answer.

OpenStudy (apoorvk):

multiply and check, what else?

OpenStudy (anonymous):

for whats question

Parth (parthkohli):

An interesting property. \(\Large \color{MidnightBlue}{\Rightarrow a^n * a^m = a^{n + m} }\)

OpenStudy (anonymous):

thx

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3^5 * 3^2 = 3^7 }\)